A train 200 m long passes a bridge 300m long. It enters the bridge with a speed of 30 M per second and leaves it with a speed of 50 M per second. What is the time taken to cross the bridge?

Grade:9

1 Answers

swastik
6 years ago
Acceleration = v-u/tA=20/tAs we know, distance = ut +1/2 x a x t^2200+300 = 30t + 1/2 x 20/t x t^2500 = 40tT= 12.5 seconds